Opened 10 years ago

Closed 8 years ago

Last modified 2 years ago

#262 closed flight processing (fixed)

IPY09/11, flight day 218/2009, Helheim

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ipy/IPY09_11-2009_218_Helheim

Data arrived from ARSF via SATA disk Daphne03 on 24/08/09.

Scientific objective: Ice thickness measurements to add to a record of two major glaciers and to create a high res DEM.

Alpha 4-High

PI: T. Murray

Sensors:

  • Eagle (08/04/2010)
  • Leica LIDAR (22/03/2010)
  • Digital Photography (11/02/2010)

Change History (19)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by knpa

Corrected logsheet project code from GB09_21

comment:3 Changed 10 years ago by benj

For day 215 I got much better IPAS results by downloading data from the hel2 station than by using Kulu. I suggest whoever processes the IPAS (+Applanix) for this does the same.

comment:4 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

Starting LIDAR processing

comment:5 Changed 9 years ago by mark1

  • Description modified (diff)

Photographs dispatched to PI on 11/02/2010

comment:6 Changed 9 years ago by emca

LiDAR processing complete. Ready for delivery check at
~airborne/workspace/IPY09_11-2009_218_Helheim/delivery/20100217/IPY09_11

comment:7 Changed 9 years ago by knpa

Delivery check completed.
All looks good apart from:

  • Need 'DEM might not be accurate' Readme in DEM dir
  • Anomalous low z values from per line statistics in Readme: -131m and -428m, is this correct?

comment:8 Changed 9 years ago by emca

LIDAR data sent to PI 22/03/2010

comment:9 Changed 9 years ago by emca

Hyperspectral data ready for delivery check at
~airborne/workspace/IPY09_11-2009_218_Helheim/delivery/20100323/IPY09_11

comment:10 Changed 9 years ago by mark1

Hyperspectral delivery checked and looks good. Ready to deliver

comment:11 Changed 9 years ago by emca

SCT values for Eagle data

Line No. SCT
1 -0.01
2 -0.05

comment:12 Changed 9 years ago by mark1

  • Description modified (diff)

Hyperspectral data dispatched to PI on (08/04/2010)

comment:13 Changed 9 years ago by emca

Camera, Hyperspectral and LIDAR data rsync-d to
/users/rsg/arsf/arsf_data/2009/flight_data/ipy/IPY09_11-2009_218_Helheim

Ready for archiving

comment:14 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

comment:15 Changed 9 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from accepted to assigned

Started archiving.

comment:16 Changed 9 years ago by knpa

Tarball created, awaiting NEODC download.

comment:17 Changed 9 years ago by knpa

Backed-up by NEODC.

Moved to archive disk.

comment:18 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:19 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.